OpenStudy (anonymous):

how do you find the range of: 4cosx+1

OpenStudy (e.mccormick):

Well, what would the range values of \(\cos x\) normally be? Then, how do those get modified by the transformations on what you have been given?

OpenStudy (anonymous):

so you raise the amplitude to 4 then shift the graph up one?

OpenStudy (anonymous):

making it go from [-1,1] to [-3,5]

OpenStudy (e.mccormick):

Yep. And look at range (output, y) values.

OpenStudy (e.mccormick):

Yah. Looks good.

OpenStudy (anonymous):

i thought if its +1 then it shift down? or is that just if your shifting it left and right then its opposite of the value

OpenStudy (ranga):

+1 shifts up.

OpenStudy (anonymous):

would cos(x+1) shift left one?

OpenStudy (e.mccormick):

Well, is it \(4cos(x+1)\) or \(4\cos (x) +1\)... ah, you are getting it.

OpenStudy (ranga):

yes.

OpenStudy (e.mccormick):

Another way of writing the range is: \(\{y|-3\le y\le 5\}\) So make sure you use the notation your class wants.

OpenStudy (anonymous):

got it thanks alot.

OpenStudy (e.mccormick):

np. have fun.